# MonoMPC: Monocular Vision Based Navigation with Learned Collision Model and Risk-Aware Model Predictive Control

### v1

@@ IV-A Implementation Details
We implemented Alg.1 in Python using Jax [26] as the GPU-accelerated linear algebra back-end. The state cost consists of a running goal cost and has the following form:
@@
IV-A1 Data Collection Pipeline
Data Collection
We collected two datasets: one via teleoperation and another using a custom planner. Data were recorded at 10 Hz, with each sample consisting of :
  • A standardized estimated point cloud of shape $300\times 2$,
@@ IV-A1 Data Collection Pipeline
  • A control sequence $\mathbf{u}=(\mathbf{v},\boldsymbol{\omega})$ (see IV-A1),
  • Ground-truth clearance $d_{gt}$ from LiDAR (see IV-A1).
Control Sequence Generation
Each sample includes $N_{p}=50$ control sequences over a 5s horizon at $\Delta t=0.1$ s. Control sequences $\mathbf{u}=(\mathbf{v},\boldsymbol{\omega})$ are generated as $\mathbf{u}\sim\mathcal{U}([-1.0,1.0])$ with $\mathbf{v}$ clipped to $[0,1]$ m/s. To get the resulting trajectories the robot state evolves via:
$\begin{bmatrix}x_{k+1}\\ y_{k+1}\\ \psi_{k+1}\end{bmatrix}=\begin{bmatrix}x_{k}+v_{k}\cos(\psi_{k})\Delta t\\ y_{k}+v_{k}\sin(\psi_{k})\Delta t\\ \psi_{k}+\omega_{k}\Delta t\end{bmatrix}$ (12)
Ground-Truth Distance Computation
For each trajectory resulting from the $j\text{-th}$ control sequence $\mathbf{u}_{j}=(\mathbf{v}_{j},\boldsymbol{\omega}_{j}),\ ,\forall j=1,\ldots,N_{p}$, the minimum distance to the LiDAR point cloud $\mathcal{P}_{\text{LiDAR}}$ is:
@@
IV-A2 Benchmarking Environment
We evaluate our approach in a model predictive control (MPC) framework using a Clearpath Jackal robot equipped with an Intel RealSense D435i RGB-D camera (RGB-only input, 69° horizontal FOV). For visualization and analysis, LiDAR data was recorded from a SICK TIM551 2D sensor. The system ran on ROS Noetic with an NVIDIA RTX 3080 GPU laptop connected to the Jackal over a local network. Depth estimation was performed in real time using TensorRT-accelerated Depth-Anything-V2 [1]. Experiments were conducted in indoor environments with static cardboard boxes to simulate clutter.
IV-B Validation on a simple example
Fig.3 shows a simple scenario where the robot needs to avoid just a single static obstacle. As can be seen, the estimated point-cloud (green) is at a large offset from the ground-truth (red) provided by the LiDAR. In spite of this, our learned collision model correctly predicts the worst-case obstacle clearance for different control samples and successfully guides the robot away from the obstacle (Fig.3(a)-(d)). Moreover, Fig.3(c)-(d) shows that our MPC framework is able to slow down the robot near the obstacles to minimize risk but compensates goal-reaching by generating higher speeds in free-space.
@@ IV-C Comparison with ROS Navigation Stack (ROSNAV)
In this subsection, we compare our method with classical planning stacks that use estimated point clouds to build costmaps. Specifically, we use navfn [^2] https://wiki.ros.org/navfn as the global planner and Dynamic Window Approach (DWA)[27] as the local planner from ROSNAV. To ensure fairness with our planner (Alg.1), which constrains linear velocities to $[0,1]$ m/s, we evaluate ROSNAV under two speed limits: 0.5 m/s and 1.0 m/s.
We observed two main failure modes with ROSNAV, both arising from inaccuracies in the estimated depth. First, due to a significant and stochastic offset between the estimated and ground-truth point clouds, dependent on obstacle placement and camera field-of-view, the robot often misidentifies free space (Fig. LABEL:fig:teaser(a)(3-4)), resulting in collisions or getting stuck. Second, as shown in our accompanying video, the goal sometimes lies within inflated costmaps derived from noisy estimates, preventing any feasible path from being generated. These issues stem from the non-trivial, time-varying nature of the estimation error, which cannot be resolved with simple heuristics.
Table I quantifies these issues: ROSNAV suffers frequent collisions, particularly at higher speeds, and fails to generate feasible plans in a substantial number of trials. In contrast, our MMD-based planner achieves up to 7x fewer collisions and consistently avoids getting stuck.
TABLE I: Comparison with ROS Navigation Stack
(60 experiments)
@@
IV-D Comparison with End-to-End Model NoMaD[5]
Figure 4: The top row shows the camera-feed when running NoMaD and our pipeline. The figure illustrates that the trajectory generated by NoMaD in (a) attempts to avoid the obstacle (red point cloud) but fails to execute a precise collision avoidance maneuver. In contrast, our method in (b) successfully deviates from the obstacle, despite the estimated point cloud (green) being offset from the ground truth. This is attributed to our learned collision model’s ability to account for uncertainty in the estimated point cloud and accurately predict the worst-case obstacle clearance.
NoMaD [5] is a foundational model for image-based goal-directed navigation. To focus on its obstacle avoidance capabilities, we evaluate it in simplified scenarios where the goal lies directly ahead and is visible from the start position, requiring only obstacle avoidance. For fair comparison, we finetune NoMaD on data from environments similar to those in Fig. LABEL:fig:teaser.
Fig.4 shows a qualitative comparison. While NoMaD trajectories tend to steer away from obstacles, they often lack the precision needed for safe navigation. In contrast, our method explicitly accounts for collision risk, resulting in safer behavior. TableII shows that NoMaD completes very few runs without collision, consistent with prior results [4], which report over $50\%$ collision rates even in simpler hallway environments. Our settings were more complex, leading to even higher failure rates.
IV-E Ablations
Table III shows that our task-aware model (Fig.2) achieves nearly 4× fewer collisions than the baseline (Fig.1) in the same test environment (Fig.LABEL:fig:teaser). This improvement stems from better variance regularization—both models predict similar means, but our model’s lower variance enables safer, less conservative plans (Fig.5). The baseline’s higher variance often causes free space to be misclassified as high-risk, making MPC planning infeasible. Simply penalizing variance in the loss (7) leads to variance collapse and overconfidence; with near-zero variance, the model becomes effectively deterministic and fails to capture risk (Table III, row 3). In contrast, our task-aware approach learns a variance that is well aligned with safe navigation.
Figure 5: The figure shows violin plots comparing the predicted variance distributions of the baseline and augmented models. The baseline model, trained only with an NLL loss, exhibits a broader spread—indicating overestimation and poor regularization of variance. In contrast, the augmented model yields a tighter distribution, reflecting task-aware variance regularization achieved through downstream risk supervision.
TABLE II: Comparison with NoMaD (60 experiments)
@@ IV-E Ablations
TABLE III: Ablations (60 experiments)
Method % Collisions % Stuck Avg. Speed (m/s) Max. Speed (m/s)
Alg.1 with $r_{MMD}^{emp}\text{-based}$
Augmented Model (Fig.2)
6.6 0 0.32 1.02
Alg.1 with
Baseline Model(Fig.1)
28.3 0 0.35 0.94
DET
41.5 0 0.52 1.01
@@ V Conclusions and Future Work
Monocular RGB-based navigation is challenging due to unreliable depth, which hinders safe collision-checking. While vision-based depth offers structural cues, its noise limits direct use. We propose a risk-aware framework that uses estimated depth as input to a probabilistic collision model, predicting worst-case obstacle clearance distributions. To regularize uncertainty, we jointly train the model and risk metric on safe and unsafe trajectories. At test time, an MPC planner uses these predictions to generate safe, goal-directed actions. Real-world experiments show our approach outperforms strong baselines like NoMaD and the ROS Navigation Stack using only monocular RGB input.
While demonstrated on a 2D ground robot, our approach readily extends to higher-dimensional platforms like quadrotors. Future directions include handling dynamic environments with time-varying collision models, accelerating planning via imitation learning, and exploring alternatives for capturing epistemic uncertainty beyond computationally costly model ensembling. We also plan to evaluate our collision model with other risk-aware MPC frameworks, such as those based on minimizing Conditional Value at Risk or enforcing chance-constraints.

### v2

@@ IV-A Implementation Details
We implemented Alg. 1 in Python using Jax [jax2018github] as the GPU-accelerated linear algebra back-end. The state cost consists of a running goal cost and has the following form:
@@
Data Collection: We collected two datasets: one via teleoperation and another using Alg. 1-first with ground-truth LiDAR for collision avoidance to train the initial collision model and subsequently using Alg. 1 with the learned collision model to collect further refinement data. Data were recorded at 10 Hz, with each sample consisting of:
  • A standardized estimated point cloud of shape $300\times 2$ obtained from Depth-Anything-V2 [NEURIPS2024_26cfdcd8],
@@
  • A control sequence $\mathbf{u}=(\mathbf{v},\boldsymbol{\omega})$ where $\mathbf{v},\boldsymbol{\omega}$ are the concatenation of the linear and angular velocities, respectively, at different timesteps,
  • Ground-truth clearance $d_{gt}$ from LiDAR (see IV-A). This is the privileged information used during training.
Control Sequence Generation: Each control sequence consists of 50 steps over a $T=5$ s horizon with sampling time of $0.1$ s. Control sequences $\mathbf{u}=(\mathbf{v},\boldsymbol{\omega})$ are generated via a uniform distribution in [1.0,1.0][-1.0,1.0] with $\mathbf{v}$ clipped to $[0,1]$ m/s. The robot state evolves via:
$x_{k+1}=x_{k}+v_{k}\cos(\psi_{k})\Delta t,\ y_{k+1}=$ $y_{k}+v_{k}\sin(\psi_{k})\Delta t$
$\psi_{k+1}=\psi_{k}+\omega_{k}\Delta t$ (10)
Ground-Truth Distance Computation For each trajectory resulting from the $j\text{-th}$ control sequence $\mathbf{u}_{j}=(\mathbf{v}_{j},\boldsymbol{\omega}_{j}),\ ,\forall j=1,\ldots,N_{p}$, the minimum distance to the LiDAR point cloud $\mathcal{P}_{\text{LiDAR}}$ is:
@@
Training Set-up: The probabilistic collision model (Sec. III-E) is trained in Equinox [kidger2021equinox] on an RTX 5090 desktop. Inputs include estimated depth, vehicle states, and control sequences. The outputs are the mean and variance of obstacle clearance distribution. We collected 80k point-cloud instances. In each of them, we used $N_{p}=50$ control sequences with randomized robot initial state to generate trajectories along which we computed the ground-truth worst-case obstacle clearance. This yielded a total of $\approx 6$M training points(point-cloud times $N_{p}$ times random initial states). The total training time was $\approx$2.5 h, following the procedure in Secs. III-DIII-E.
Benchmarking Setup: We evaluate our approach using a Clearpath Jackal robot equipped with an Intel RealSense D435i RGB-D camera (RGB-only input, 69° horizontal FOV). For visualization and analysis, LiDAR data was recorded from a SICK TIM551 2D sensor. The system ran on ROS Noetic with an NVIDIA RTX 3080 GPU laptop connected to the Jackal over a local network. Depth estimation was performed in real time using TensorRT-accelerated Depth-Anything-V2 [NEURIPS2024_26cfdcd8]. Experiments were conducted in indoor environments with static cardboard boxes to simulate clutter.
Figure 3: Box plot of minimum residuals between sampled worst-case obstacle clearances and ground truth. The augmented model achieves lower median error and reduced variability compared to the baseline, indicating sharper and more consistent alignment with ground-truth worst-case obstacle clearances.
Figure 4: Comparison between our approach (blue) and MonoNav [simon2023mononav] (red); goal in green. The noise in the estimated depth translates to erroneous 3D occupancy maps-the yellow cuboids (ground-truth obstacles) do not align with their reconstructed point clouds resulting in MonoNav getting stuck or colliding (b-c). In contrast, our approach is able to avoid the yellow cuboids based on the noisy estimated depth. We do not use MonoNav’s 3D reconstruction; trajectories are overlaid solely for visualization and comparison.
IV-B Validation of the Learned Collision Model
In this subsection, we validate the task-aware learning of the collision model (Fig. 2) and compare with the baseline approach (Fig. 1).
Predictive Accuracy:We analyze residuals between sampled worst-case obstacle clearances and ground truth. For each test case, we record the minimum absolute residual between samples from the predictive distribution and the ground truth, indicating whether any sample captures the true worst-case clearance. As shown in Fig. 3, the task-aware augmented model of Fig. 2 yields lower median residuals and a tighter inter-quartile range than the baseline (Fig. 1), indicating more consistent and reliable predictions.
Distributional Metrics: We evaluate three distributional metrics - Energy Score (ES), Probability Integral Transform (PIT), and Expected Calibration Error (ECE). The ES [Gneiting01032007] measures both accuracy and sharpness, with lower values indicating predictions closer to the ground truth. Perfect calibration yields uniformly distributed PIT values [repec:nbr:nberte:0215] on $[0,1]$, while deviations from uniformity are summarized by ECE [Naeini2015ObtainingWC], where $\text{ECE}\approx 0$ denotes ideal calibration.
Additionally, to evaluate safety, we compute four outcomes on the test-set: Correctly Safe (CS), Correctly Unsafe (CU), Unsafe (U), and Conservative (C). We define safety-critical rates of Unsafe (UR) and Conservative (CR):
$\text{UR}=\text{U}/(\text{U}+\text{CU})\times 100\%,\quad\text{CR}=\text{C}/(\text{C}+\text{CS})\times 100\%$
UR measures dangerous failures where unsafe states are overlooked, while CR captures over-conservatism where safe states are misclassified. For each outcome, risk $r_{MMD}^{emp}$ is computed from predicted clearances and compared against the ground-truth risk from $d_{gt}$.
Table I shows that the $r_{MMD}^{emp}$-based augmented collision model (Fig. 2) surpasses the baseline (Fig. 1) across all metrics: UR $\downarrow$50%, CR $\downarrow$80%, ES $\downarrow$74%, and ECE $\downarrow$20%. These gains stem from task-aware variance supervision aligning uncertainty with collision risk, unlike the NLL-only baseline(Fig. 1), which produces mis-calibrated and unreliable predictions. The same trend appears in navigation results (Table II), where the task-aware model reduces collisions by $4$ times. Table II also shows performance with deterministic collision model that performs the worst.
@@ IV-C Comparison with ROS Navigation Stack (ROSNAV)
We benchmark our method against a classical planning stack from ROSNAV, which uses global planning followed by Dynamic Window Approach (DWA) for local control on costmaps built from estimated point clouds. To ensure a fair comparison with our planner’s constraints (Alg. 1), we evaluate ROSNAV at two speed limits: 0.5 m/s and 1.0 m/s. This classical approach fails mainly due to its inability to handle noise in the estimated depth. Offsets between estimated and true point clouds lead to false free space and collisions (Fig. LABEL:fig:teaser(a)(3-4)), while noisy estimates inflate costmaps near goals, causing planning failures. These results show that such complex, dynamic errors cannot be corrected by simple heuristics. Table III quantifies these trends: ROSNAV often collides or fails at higher speeds, whereas our $r_{MMD}^{emp}$-based planner cuts collisions by up to 7 times and eliminates planning failures.
TABLE I: Validation Metrics for the Learned Collision Model
Method
UR (%)
(lower is better)
CR (%)
(lower is better)
ES
(lower is better)
ECE over PIT
(lower is better)
Alg. 1 with $r_{MMD}^{emp}\text{-based}$
Augmented Model (Fig. 2)
1.3 5.36 0.0236 0.0893
Alg. 1 with
Baseline Model (Fig. 1)
2.97 30.88 0.0917 0.1142
TABLE II: Navigation with and without Task-Aware Training
Method % Collisions % Stuck Avg. Speed (m/s) Max. Speed (m/s)
Alg. 1 with $r_{MMD}^{emp}\text{-based}$
Augmented Model (Fig. 2)
6.6 0 0.32 1.02
Alg. 1 with
Baseline Model(Fig. 1)
28.3 0 0.35 0.94
Deteministic
41.5 0 0.52 1.01
TABLE III: Comparison with ROS Navigation Stack
@@
IV-D Comparison with MonoNav [simon2023mononav]
To highlight the benefits of our uncertainty-aware framework, we benchmark it against MonoNav [simon2023mononav], a baseline that plans directly on 3D reconstructions from estimated depth. In a series of quadcopter navigation experiments (40 total) through increasingly cluttered 2.5D environments (Fig. 4), we demonstrate the critical failure modes of ignoring depth uncertainty.
As shown in Fig. 4, noise in MonoNav’s depth perception causes flawed reconstructions-tolerable in sparse scenes (a) but fatal in cluttered ones, leading to entrapment (b) or collision (c). Its simplistic primitive-based planner further limits adaptability to such errors. In contrast, our method explicitly incorporates depth uncertainty into risk-aware trajectory optimization, enabling successful navigation even in complex scenes. The executed trajectory (blue) consistently avoids all ground-truth obstacles (yellow cuboids), as confirmed by Table IV and the accompanying video showing smoother, faster motions than MonoNav.
TABLE IV: Comparison with MonoNav
Method Easy Setting     Moderately Difficult Setting
% Collisions % Stuck     % Collisions % Stuck
Alg. 1 with $r_{MMD}^{emp}\text{-based}$
Augmented Model
0 0     0 0
MonoNav [simon2023mononav]
0 0     20 80
IV-E Comparison with End-to-End Model NoMaD[10610665]
NoMaD [10610665] is a foundational model for image-based goal-directed navigation. To focus on its obstacle avoidance capabilities, we evaluate it in simplified scenarios where the goal lies directly ahead and is visible from the start position, requiring only obstacle avoidance. We fine-tuned NoMaD using tele-operation data ($\sim$10k training points) and evaluated it in a test environment very similar to its training setup, giving it the best possible conditions to perform reliably.
Table V shows that NoMaD completes very few runs without collision, consistent with prior results [simon2023mononav], which report over $50\%$ collision rates even in simpler hallway environments. Our settings were more complex, leading to even higher failure rates.
TABLE V: Comparison with NoMaD
@@ IV-E Comparison with End-to-End Model NoMaD[10610665]
IV-F Computation Time
On an RTX 3080 laptop, planning runs in $0.01$ s with $0.02$ s for parallel depth estimation. On a Jetson Orin, planning time is $0.065$ s, thus demonstrating the feasibility of deploying our system on mobile robots with real-time re-planning.
@@ V Conclusions and Future Work
In this work, we showed that depth estimates from vision foundation models are unreliable for direct use in autonomous navigation. For example, occupancy maps built from them are error-prone, causing failures in cluttered scenes (e.g., ROS Navigation Stack, MonoNav[simon2023mononav]). On the other hand, end-to-end vision-based methods such as NoMaD [10610665] lack explicit constraint handling and also perform poorly. We address these limitations by formulating monocular vision-based navigation as a risk-aware planning problem over a learned probabilistic collision model. Our training pipeline enforces well-calibrated uncertainty, improving downstream risk estimation and planning performance. The proposed approach operates in real-time on both commodity laptops and Jetson Orin, with successful demonstrations on ground and aerial platforms.
A limitation of our approach is the lack of temporal memory, as it ignores past images or point-clouds. This can cause oscillatory behavior in cluttered environments. Future work will focus on incorporating temporal information to mitigate this issue. Future work also includes extending our collision model to dynamic environments with time-varying collision models and accelerating planning through imitation learning. We also plan to explore end-to-end frameworks where collision model and action policy are jointly trained through demonstration data.